OpenStudy (anonymous):

Given the equation y − 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

y = mx + c

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

3x − 4y = −40

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y = 3/4x + 12

OpenStudy (anonymous):

y − 4 = 3/4(x + 8) y = 3/4(x + 8) + 4 y = (3/4)x + 6 + 4 y = (3/4)x + 10

OpenStudy (anonymous):

thank you